can you help me with this question?
(4x) (6,3x) M= 1/2

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(31513)

What is the question?

Answer by stanbon(57959)

(4x) (6.3x) M= 1/2
---
25.2x^2 = 1/2
---
x^2 = 0.5/25.2
---
x = +/- sqrt[0.5/25.2]
---
x = +/- 0.1409
=====
Cheers,
Stan H.